OpenStudy (anonymous):

How do i solve (x+6)^2/3=9?

OpenStudy (anonymous):

multiply both sides by 3 (x+6)^2= 27

OpenStudy (anonymous):

simplify the square x^2+12x+36=27

OpenStudy (anonymous):

substract 27 from both the sides x^2+12x+9

OpenStudy (anonymous):

use the equation \[x= (-b \pm \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

where a=1 b=12 c=9

OpenStudy (mertsj):

\[(x+6)^2=27\] \[\sqrt{(x+6)^2}=\pm \sqrt{27}\] \[x+6=\pm3\sqrt{3}\]

OpenStudy (mertsj):

\[x=-6\pm3\sqrt{3}\]

OpenStudy (anonymous):

@d3v13 would it be (x + 6)^(2/3) = 9 (x + 6) = 9^(3/2) x + 6 = -27 , 27 x = -33 , 21

OpenStudy (mertsj):

\[(x+6)^{\frac{2}{3}}=9\] \[((x+6)^{\frac{2}{3}})^{\frac{3}{2}}=9^{\frac{3}{2}}\] \[x+6=27\] \[x=21\]

OpenStudy (anonymous):

^is right. 9^(1/2)=3 3^3=27

OpenStudy (anonymous):

and the reason behind that is sqrt(-3^2) is not equal to sqrt(9)